OpenStudy (anonymous):

@jim_thompson5910 Q#1 :The price of a bag of oranges is directly proportional to the number of oranges packaged in the bag. The table shows this relationship. What is the constant of proportionality? Oranges (x) 8 12 16 Price ($) (y) 2 3 4 my answer :4

jimthompson5910 (jim_thompson5910):

y = k*x 2 = k*8 k = ???

OpenStudy (anonymous):

i got 18

OpenStudy (anonymous):

is it right

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

2 = k*8 divide both sides by 8

OpenStudy (anonymous):

very confused this question is very hard

jimthompson5910 (jim_thompson5910):

what is 2/8 equal to?

OpenStudy (anonymous):

0.25

jimthompson5910 (jim_thompson5910):

yes or 1/4

OpenStudy (anonymous):

Q#1: Madison bought a gift for $17.55. Sales tax of 8% was included in the price she paid. What was the price of the gift before the tax was added? A. $15.75 B. $16.15 C. $16.25 D. $16.47 my answer : C

jimthompson5910 (jim_thompson5910):

yeah you solve 1.08*x = 17.55 for x to get x = 16.25

OpenStudy (anonymous):

that's what i did

OpenStudy (anonymous):

Q#3:A $6000 investment earns $690 simple interest in two years. What is the annual interest rate? A. 1.15% B. 5.75% C. 11.5% D. 57.5% my answer:B

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Q#4 :Alan estimated that he would score 140 points the next time he bowled. He actually scored 152 points. What was the percent error in Alan’s estimate? Round to the nearest tenth of a percent, if needed. A. 6% B. 7.9% C. 8.6% D. 12% my answer :B